# Revisiting Non-Rotating Star Models: Classical Existence and Uniqueness Theory and Scaling Relations

### v1

@@ 2.2 Existence Results for Non-rotating Bodies
$\int\left(\sigma_{R}\right)^{\frac{4}{3}}$ $=\int_{\left\{x\mid\sigma_{R}(x)<s_{0}\right\}}\left(\sigma_{R}\right)^{\frac{1}{3}}\cdot\sigma_{R}+\int_{\left\{x\mid\sigma_{R}(x)\geq s_{0}\right\}}\left(\sigma_{R}\right)^{\frac{4}{3}}$
$\leq s_{0}^{\frac{1}{3}}\cdot m+\frac{1}{(\widetilde{C}+2)}\int A\left(\sigma_{R}\right)$
@@
Recall we have $E_{0}^{\prime}\left(\sigma_{m}\right)(x)=A^{\prime}\left(\sigma_{m}(x)\right)-V_{\sigma_{m}}(x)$ by Lemma 2.8, then we can also write
@@
In fact, we first apply (generalized) Lagrange multiplier theorem Theorem A.1 to obtain (22)(23)(24) holds true a.e. in the region described above, then we get
@@
Although the sign of $\lambda_{R}$ is currently unknown (it will be shown negative later), we can still show $\sigma_{R}{}^{\frac{1}{3}}\leq cV_{\sigma_{R}}$ a.e. on the set where $\sigma_{R}\geq K_{R}$ for some $K_{R}>0$ by absorbing $\lambda_{R}$ when $\sigma_{R}$ is large enough [1, Lemma 3]. Similar to the previous arguments in [7, Theorem 2.25 (iii-vi)] and Remark 2.16, we can show $\sigma_{R}$ is continuous in $B_{R}(0)$ and thereby replace almost everywhere (a.e.) by everywhere.
One may consider use similar arguments as above (bootstrap method) to show Lemma 2.22. But we need to realize that this time $\sigma_{R}^{\frac{1}{3}}\leq cV_{\sigma_{R}}$ a.e. on the set where $\sigma_{R}\geq K_{R}$, here $K_{R}$ depends on $\lambda_{R}$ thus may not be a uniform constant.
However, this problem can be solved. We claim $K_{R}$ is indeed a uniform constant for sufficient large $R$. We can show $\lambda_{R}<0$ for sufficient large $R$. To show $\lambda_{R}<0$, we choose suitable region $U_{R}$ where (22) holds that $\lambda_{R}\leq E_{0}{}^{\prime}\left(\sigma_{R}\right)$, analyze the behaviors of terms in $E_{0}{}^{\prime}\left(\sigma_{R}\right)$ as $R\rightarrow\infty$, and obtain $\lambda_{R}\leq o\left(R^{-1}\right)-cmR^{-1}<0$ on $U_{R}$. For example, if $x\in B_{R}(0)$, since $\sigma_{R}=0$ outside $B_{R}(0)$, we have
$V_{\sigma_{R}}(x)=\int\frac{\sigma_{R}(y)}{|x-y|}d^{3}y=\int_{B_{R}(0)}\frac{\sigma_{R}(y)}{|x-y|}d^{3}y\geq\frac{1}{2R}\int_{B_{R}(0)}\sigma_{R}(y)d^{3}y=\frac{m}{2R}$ (26)
which gives us the last term $cmR^{-1}$. One can also see how to get the term $o(R^{-1})$ in [1, Lemma 5]. After showing $\lambda_{R}<0$, thanks to (24) we have $A^{\prime}\left(\sigma_{R}\right)\leq V_{\sigma_{R}}$ for $R$ large enough. Recall Lemma 2.2 (iii) tells us $\sigma_{R}^{\frac{1}{3}}\leq A^{\prime}(\sigma_{R})$ when $\sigma_{R}>N$ for some $N>0$. It implies we can take $K_{R}=N$ such that it is a uniform constant for $R\geq R_{1}$, where $R_{1}$ is sufficiently large. Therefore, we can truly apply bootstrap method to prove $\exists C_{1}>0$, such that $\left\|\sigma_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C_{1}$ for $R\geq R_{1}$. Take $C=\max\{C_{1},R_{1}\}$, we know $\left\|\sigma_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C$ for $R>0$. This finishes the proof of Lemma 2.22 (Step (1)). ∎
@@
(F3) is important to show the uniform bound of $\sigma_{R}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$. It helps to show $\int\left(\sigma_{R}\right)^{\frac{4}{3}}\leq s_{0}^{\frac{1}{3}}\cdot m+\frac{1}{(\widetilde{C}+2)}\int A\left(\sigma_{R}\right)$. If we replace $\frac{4}{3}$ by any smaller number in (F3), we could not obtain such uniform bound since we would not have $(\widetilde{C}+1)\int\left(\sigma_{R}\right)^{\frac{4}{3}}\leq E_{0}(\widetilde{\rho})+\widetilde{C}\int\left(\sigma_{R}\right)^{\frac{4}{3}}$ as above, as Proposition 2.17 suggests.
@@
$\exists\widetilde{R}_{0}$, such that $\forall R>\widetilde{R}_{0},E_{0}\left(\sigma_{R}\right)\leq a<0$ and $\lambda_{R}\leq l<0$, where $\lambda_{R}$ is the corresponding Lagrange multiplier.
@@
For energies, we choose a text function $\rho_{R}\in W_{R}$ and estimate $E\left(\rho_{R}\right)=o\left(R^{-1}\right)-cR^{-1}$ as $R\rightarrow\infty$, see [1, Lemma 6], then $\exists R_{2}$ such that $\forall R\geq R_{2},E\left(\sigma_{R}\right)\leq E\left(\sigma_{R_{2}}\right)\leq E\left(\rho_{R_{2}}\right):=a<0$.
@@
$a\leq-E\left(\sigma_{R}\right)\leq\frac{1}{2}\int\sigma_{R}V_{\sigma_{R}}\leq\frac{1}{2}m\left\|V_{\sigma_{R}}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)},$
then for $R$ large enough, $\left\|V_{\sigma_{R}}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\geq-\frac{2a}{m}$, which can help to imply $\sigma_{R}$ cannot become uniformly diffuse as $R\rightarrow\infty$. That is, we claim $\exists\epsilon_{0}$, such that $\forall R>0,\exists x_{R}\in\mathbb{R}^{3}$, we have
$\int_{\left|y-x_{R}\right|<1}\sigma_{R}(y)d^{3}y\geq\epsilon_{0}$
In fact, let $\epsilon_{R}=\sup\limits_{x}\int_{|y-x|<1}\rho_{R}$, and estimate $V_{\sigma_{R}}(x)$ as
$V_{\sigma_{R}}(x)$ $=\int_{|y-x|<1}\frac{\sigma_{R}(y)}{|y-x|}d^{3}y+\int_{1<|y-x|<r}\frac{\sigma_{R}(y)}{|y-x|}d^{3}y+\int_{|y-x|>r}\frac{\sigma_{R}(y)}{|y-x|}d^{3}y$
@@
It’s easy to see $V_{3}\leq\frac{M}{r}$. Since the shell 1<|yx|<r1<|y-x|<r can be covered by finite balls of radius 1 (it’s 3-dimensional space thus the number of balls can be $\leq cr^{3}$), $V_{2}\leq c\epsilon_{R}r^{3}$. Thanks to Lemma 2.22 $\left(\left\|\sigma_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq k_{1}\right)$ and Proposition A.5, we have $V_{1}\leq c\left(\epsilon_{R}^{a}+\epsilon_{R}^{b}\right)$ where $a>0,b>0$. Let $g(r)=\frac{M}{r}+c\epsilon_{R}r^{3}+c\left(\epsilon_{R}^{a}+\epsilon_{R}^{b}\right)$, where $r>0$. We know it has minimal value $g_{\min}=\widetilde{c}\epsilon_{R}^{\frac{1}{4}}+c\left(\epsilon_{R}^{a}+\epsilon_{R}^{b}\right)$ at $r_{\min}=\left(\frac{M}{3c\epsilon_{R}}\right)^{\frac{1}{4}}$. If $r_{\min}\geq 1$, then $\left\|V_{\sigma_{R}}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq\widetilde{c}\epsilon_{R}^{\frac{1}{4}}+c\left(\epsilon_{R}^{a}+\epsilon_{R}^{b}\right)$. On the other hand, $\left\|V_{\sigma_{R}}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\geq-\frac{2e}{m}$. Therefore, $\exists\widetilde{\epsilon}>0,\widetilde{R}>0$, such that $\forall R>\widetilde{R}$, we have $\epsilon_{R}>\widetilde{\epsilon}$. If $r_{\text{min }}<1$, then $\epsilon_{R}>\frac{M}{3c}$, with an abuse of notation, $\epsilon_{R}>\widetilde{\epsilon}$ still holds by letting $\widetilde{\epsilon}=\min\left\{\widetilde{\epsilon},\frac{M}{3c}\right\}$. If $R\leq\widetilde{R},\sigma_{R}\in W_{\widetilde{R}}$, thus spt $\sigma_{R}$ is covered by finite balls with radius 1 (the number of balls $C$ is independent of $R$). Since the mass of $\sigma_{R}$ is $m$, we know $\epsilon_{R}>\frac{m}{2C}$. Let $\epsilon=\min\left\{\widetilde{\epsilon},\frac{m}{2C}\right\}$, we know $\epsilon_{R}>\epsilon$ for all $R$, which proves the claim.
By the construction of $W_{R}$ we get $\sigma_{R}$ is axially symmetric, then $x_{R}$ mentioned in the claim should satisfies $r\left(x_{R}\right)\leq r_{0}$.
@@
The result that the energy is negative is nice, since later we will show $\sigma_{R}$ is truly a minimizer on $W_{\infty}=mR{(\mathbb{R}^{3})}$ thus $E_{0}\left(\sigma_{R}\right)$ is the minimal energy $e_{0}(m)$, which is less than 0.
@@
We first show $V_{\sigma_{R}}(x)\leq-l$ holds when the distance from $x$ to the z-axis $r(x)$ and $R$ is large enough, then due to (24) and Lemma 2.25 we know $\sigma_{R}(x)=0$. Let $f(\pi)=c_{1}\left(\pi^{-a}+\pi^{-b}\right)+c_{2}\pi^{-1}r^{3}+c_{3}r^{-1}$, where $a>0,b>0$, if we fix $r=r_{0}:=1+\frac{-3c_{3}}{l}$, then $c_{3}r^{-1}<-\frac{l}{3}$, and we have $f(\pi)\leq-l$ when $\pi$ is large enough.
Now we claim $\exists\pi_{0}>2r_{0}-1,\forall\pi>\pi_{0},\forall x$ with $r(x)>\pi,\forall r\in\left(1,\frac{1+\pi}{2}\right)$ we have $V_{\sigma_{R}}(x)\leq f(\pi)\leq-l$.
In fact, let $\epsilon_{\pi,R}=\sup\limits_{r(x)>\frac{\pi}{2}}\int_{|y-x|<1}\sigma_{R}(y)d^{3}y$, then $\exists x_{\pi,R}\in\mathbb{R}^{3}$ with $r\left(x_{\pi,R}\right)>\frac{\pi}{2}$, such that
$\int_{\left|y-x_{\pi,R}\right|<1}\sigma_{R}(y)d^{3}y>\epsilon_{\pi,R}-\frac{1}{\pi}$
On the other hand, due to axial symmetry of $\sigma_{R}(y),\sigma_{R}(y)$ has mass $\geq{cr}\left(x_{\pi,R}\right)$ $\left(\epsilon_{\pi,R}-\frac{1}{\pi}\right)$ in the torus obtained by revolving the ball $\left|y-x_{\pi,R}\right|<1$ around the z-axis, which implies $\epsilon_{\pi,R}\leq\frac{m}{{cr}\left(x_{\pi,R}\right)}+\frac{1}{\pi}<\frac{2m}{c\pi}+\frac{1}{\pi}$. With an abuse of notation, we have $\epsilon_{\pi,R}<\frac{c}{\pi}$.
$\forall x$ with $r(x)>\pi$, notice due to $r<\frac{1+\pi}{2}$, we can choose finite balls centered at $x_{n}$ of radius 1 with $r\left(x_{n}\right)>\frac{\pi}{2}$ covering 1<|yx|<r1<|y-x|<r, and the number of balls can be $\leq cr^{3}$. Similar as Lemma 2.25, together with $\epsilon_{\pi,R}<\frac{c}{\pi}$, we have
$V_{\sigma_{R}}(x)$ $=\int_{|y-x|<1}\frac{\sigma_{R}(y)}{|y-x|}d^{3}y+\int_{1<|y-x|<r}\frac{\sigma_{R}(y)}{|y-x|}d^{3}y+\int_{|y-x|>r}\frac{\sigma_{R}(y)}{|y-x|}d^{3}y$

### v2

@@ 2.2 Existence Results for Non-rotating Bodies
$\int_{\mathbb{R}^{3}}A\left(\sigma_{R}\right)\,dx\leq E_{0}(\widetilde{\rho})+Cm^{\frac{2}{3}}\int_{\mathbb{R}^{3}}\sigma_{R}^{\frac{4}{3}}\,dx$
where $m=\int_{\mathbb{R}^{3}}\sigma_{R}\,dx$. Let $\widetilde{C}=Cm^{\frac{2}{3}}$, by Lemma 2.2 (ii), $\exists s_{0}>0$, such that $\forall s\geq s_{0}$, $(\widetilde{C}+2)s^{\frac{4}{3}}\leq A(s)$. Notice the mass of $\sigma_{R}$ is $m$. Then
$\int_{\mathbb{R}^{3}}\left(\sigma_{R}\right)^{\frac{4}{3}}\,dx$ $=\int_{\left\{x\mid\sigma_{R}(x)<s_{0}\right\}}\left(\sigma_{R}\right)^{\frac{1}{3}}\cdot\sigma_{R}\,dx+\int_{\left\{x\mid\sigma_{R}(x)\geq s_{0}\right\}}\left(\sigma_{R}\right)^{\frac{4}{3}}\,dx$
$\leq s_{0}^{\frac{1}{3}}\cdot m+\frac{1}{(\widetilde{C}+2)}\int_{\mathbb{R}^{3}}A\left(\sigma_{R}\right)\,dx$
[…]
$(\widetilde{C}+1)\int_{\mathbb{R}^{3}}\left(\sigma_{R}\right)^{\frac{4}{3}}\,dx\leq(\widetilde{C}+2)\int_{\mathbb{R}^{3}}\left(\sigma_{R}\right)^{\frac{4}{3}}\,dx-(\widetilde{C}+2)s_{0}^{\frac{1}{3}}\cdot m\leq\int_{\mathbb{R}^{3}}A\left(\sigma_{R}\right)\,dx\leq E_{0}(\widetilde{\rho})+\widetilde{C}\int_{\mathbb{R}^{3}}\left(\sigma_{R}\right)^{\frac{4}{3}}\,dx$
@@
Recall we have $E_{0}^{\prime}\left(\sigma\right)(x)=A^{\prime}\left(\sigma(x)\right)-V_{\sigma}(x)$ by Lemma 2.9, then we can also write
@@
In fact, we first apply (generalized) Lagrange multiplier theorem [17, Appendix A] to obtain (21)(22)(23) holds true a.e. in the region described above, then we get
@@
Although the sign of $\lambda_{R}$ is currently unknown (it will be shown negative later), we can still show $\sigma_{R}{}^{\frac{1}{3}}\leq cV_{\sigma_{R}}$ a.e. on the set where $\sigma_{R}\geq K_{R}$ for some $K_{R}>0$ by absorbing $\lambda_{R}$ when $\sigma_{R}$ is large enough [4, Lemma 3]. Similar to the arguments in [17, Section 2], we can show $\sigma_{R}$ is continuous in $B_{R}(0)$ and thereby replace almost everywhere (a.e.) by everywhere.
One may consider using arguments similar to those above (bootstrap method) to show Lemma 2.21. But we need to realize that this time $\sigma_{R}^{\frac{1}{3}}\leq cV_{\sigma_{R}}$ a.e. on the set where $\sigma_{R}\geq K_{R}$, here $K_{R}$ depends on $\lambda_{R}$ thus may not be a uniform constant.
However, this problem can be solved. We claim $K_{R}$ is indeed a uniform constant for sufficiently large $R$. We can show $\lambda_{R}<0$ for sufficiently large $R$. To show $\lambda_{R}<0$, we choose a suitable region $U_{R}$ where (21) holds that $\lambda_{R}\leq E_{0}{}^{\prime}\left(\sigma_{R}\right)$, analyze the behaviors of terms in $E_{0}{}^{\prime}\left(\sigma_{R}\right)$ as $R\rightarrow\infty$, and obtain $\lambda_{R}\leq o\left(R^{-1}\right)-cmR^{-1}<0$ on $U_{R}$. For example, if $x\in B_{R}(0)$, since $\sigma_{R}=0$ outside $B_{R}(0)$, we have
$V_{\sigma_{R}}(x)=\int_{\mathbb{R}^{3}}\frac{\sigma_{R}(y)}{|x-y|}\,dy=\int_{B_{R}(0)}\frac{\sigma_{R}(y)}{|x-y|}\,dy\geq\frac{1}{2R}\int_{B_{R}(0)}\sigma_{R}(y)\,dy=\frac{m}{2R}$ (25)
which gives us the last term $cmR^{-1}$. One can also see how to get the term $o(R^{-1})$ in [4, Lemma 5]. After showing $\lambda_{R}<0$, thanks to (23) we have $A^{\prime}\left(\sigma_{R}\right)\leq V_{\sigma_{R}}$ for $R$ large enough. Recall Lemma 2.2 (iii) tells us $\sigma_{R}^{\frac{1}{3}}\leq A^{\prime}(\sigma_{R})$ when $\sigma_{R}>N$ for some $N>0$. It implies we can take $K_{R}=N$ such that it is a uniform constant for $R\geq R_{1}$, where $R_{1}$ is sufficiently large. Therefore, we can truly apply bootstrap method to prove $\exists C_{1}>0$, such that $\left\|\sigma_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C_{1}$ for $R\geq R_{1}$. Take $C=\max\{C_{1},R_{1}\}$, we know $\left\|\sigma_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C$ for $R>0$. This finishes the proof of Lemma 2.21 (Step (1)). ∎
@@
(F3) is important to show the uniform bound of $\sigma_{R}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$. It helps to show $\int_{\mathbb{R}^{3}}\left(\sigma_{R}\right)^{\frac{4}{3}}\,dx\leq s_{0}^{\frac{1}{3}}\cdot m+\frac{1}{(\widetilde{C}+2)}\int_{\mathbb{R}^{3}}A\left(\sigma_{R}\right)\,dx$. If we replace $\frac{4}{3}$ by any smaller number in (F3), we could not obtain such uniform bound since we would not have $(\widetilde{C}+1)\int_{\mathbb{R}^{3}}\left(\sigma_{R}\right)^{\frac{4}{3}}\,dx\leq E_{0}(\widetilde{\rho})+\widetilde{C}\int_{\mathbb{R}^{3}}\left(\sigma_{R}\right)^{\frac{4}{3}}\,dx$ as above, as Proposition 2.16 suggests.
@@
$\exists\widetilde{R}_{0},\alpha,l$, such that $\forall R>\widetilde{R}_{0}$, $E_{0}\left(\sigma_{R}\right)\leq\alpha<0$ and $\lambda_{R}\leq l<0$, where $\sigma_{R}$ is the energy minimizer in $W_{R}$, and $\lambda_{R}$ is the corresponding Lagrange multiplier.
@@
For energies, we choose a text function $\rho_{R}(x)=R^{-3}\rho(R^{-1}x)\in W_{R}$ for some suitable function $\rho$, and estimate $E\left(\rho_{R}\right)=o\left(R^{-1}\right)-cR^{-1}$ as $R\rightarrow\infty$, see [4, Lemma 6], then $\exists R_{2}$ such that $\forall R\geq R_{2}$, $E\left(\sigma_{R}\right)\leq E\left(\sigma_{R_{2}}\right)\leq E\left(\rho_{R_{2}}\right)\coloneq\alpha<0$.
@@
$\alpha\leq-E\left(\sigma_{R}\right)\leq\frac{1}{2}\int_{\mathbb{R}^{3}}\sigma_{R}V_{\sigma_{R}}\,dx\leq\frac{1}{2}m\left\|V_{\sigma_{R}}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$
then for $R$ large enough, $\left\|V_{\sigma_{R}}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\geq-\frac{2\alpha}{m}$, which can help to imply $\sigma_{R}$ cannot become uniformly diffuse as $R\rightarrow\infty$. That is, we claim $\exists\epsilon_{0}$, such that $\forall R>0,\exists x_{R}\in\mathbb{R}^{3}$, we have
$\int_{\left|y-x_{R}\right|<1}\sigma_{R}(y)\,dy\geq\epsilon_{0}$
In fact, let $\epsilon_{R}=\sup\limits_{x}\int_{|y-x|<1}\rho_{R}\,dy$, and estimate $V_{\sigma_{R}}(x)$ as
$V_{\sigma_{R}}(x)$ $=\int_{|y-x|<1}\frac{\sigma_{R}(y)}{|y-x|}\,dy+\int_{1<|y-x|<r}\frac{\sigma_{R}(y)}{|y-x|}\,dy+\int_{|y-x|>r}\frac{\sigma_{R}(y)}{|y-x|}\,dy$
@@
It’s easy to see $V_{3}\leq\frac{M}{r}$. Since the shell 1<|yx|<r1<|y-x|<r can be covered by finite balls of radius 1 (it is 3-dimensional space thus the number of balls can be $\leq cr^{3}$), then $V_{2}\leq c\epsilon_{R}r^{3}$. Thanks to Lemma 2.21 $\left(\left\|\sigma_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq k_{1}\right)$ and Proposition A.3, we have $V_{1}\leq c\left(\epsilon_{R}^{a}+\epsilon_{R}^{b}\right)$ where $a>0,b>0$. Let $g(r)=\frac{M}{r}+c\epsilon_{R}r^{3}+c\left(\epsilon_{R}^{a}+\epsilon_{R}^{b}\right)$, where $r>0$. We know it has minimal value $g_{\min}=\widetilde{c}\epsilon_{R}^{\frac{1}{4}}+c\left(\epsilon_{R}^{a}+\epsilon_{R}^{b}\right)$ at $r_{\min}=\left(\frac{M}{3c\epsilon_{R}}\right)^{\frac{1}{4}}$. If $r_{\min}\geq 1$, then $\left\|V_{\sigma_{R}}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq\widetilde{c}\epsilon_{R}^{\frac{1}{4}}+c\left(\epsilon_{R}^{a}+\epsilon_{R}^{b}\right)$. On the other hand, $\left\|V_{\sigma_{R}}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\geq-\frac{2e}{m}$. Therefore, $\exists\widetilde{\epsilon}>0,\widetilde{R}>0$, such that $\forall R>\widetilde{R}$, we have $\epsilon_{R}>\widetilde{\epsilon}$. If $r_{\text{min }}<1$, then $\epsilon_{R}>\frac{M}{3c}$, with an abuse of notation, $\epsilon_{R}>\widetilde{\epsilon}$ still holds by letting $\widetilde{\epsilon}=\min\left\{\widetilde{\epsilon},\frac{M}{3c}\right\}$. If $R\leq\widetilde{R},\sigma_{R}\in W_{\widetilde{R}}$, thus spt $\sigma_{R}$ is covered by finite balls with radius 1 (the number of balls $C$ is independent of $R$). Since the mass of $\sigma_{R}$ is $m$, we know $\epsilon_{R}>\frac{m}{2C}$. Let $\epsilon=\min\left\{\widetilde{\epsilon},\frac{m}{2C}\right\}$, we know $\epsilon_{R}>\epsilon$ for all $R$, which proves the claim.
By the construction of $W_{R}$ we get $\sigma_{R}$ is axially symmetric, then $x_{R}$ mentioned in the claim should satisfy $r\left(x_{R}\right)\leq r_{0}$.
@@
The result that the energy is negative is nice, since later we will show $\sigma_{R}$ is truly a minimizer on $W_{\infty}=mR{(\mathbb{R}^{3})}$, then $E_{0}\left(\sigma_{R}\right)$ is the minimal energy $e_{0}(m)$, which is less than 0.
@@
We first show $V_{\sigma_{R}}(x)\leq-l$ holds when the distance from $x$ to the z-axis $r(x)$ and $R$ is large enough, then due to (23) and Lemma 2.24 we know $\sigma_{R}(x)=0$. Let $a,b>0$ be as in the estimate $V_{1}\leq c(\epsilon_{R}^{a}+\epsilon_{R}^{b})$ from the proof of Lemma 2.24, and let $c_{1},c_{2},c_{3}>0$ be sufficiently large constants. Then define
$f(r,\pi)=c_{1}(\pi^{-a}+\pi^{-b})+c_{2}\pi^{-1}r^{3}+c_{3}r^{-1}$
Notice if we choose $r\geq r_{0}:=1+\frac{-3c_{3}}{l}$, then $c_{3}r^{-1}<-\frac{l}{3}$, and we have $f(r,\pi)\leq-l$ when $\pi$ is large enough.
Now we claim $\exists\pi_{0}>2r_{0}-1$, $\forall\pi>\pi_{0}$, $\forall x$ with $r(x)>\pi$, $\forall r\in\left(r_{0},\frac{1+\pi}{2}\right)$ we have $V_{\sigma_{R}}(x)\leq f(r,\pi)\leq-l$.
In fact, let $\epsilon_{\pi,R}=\sup\limits_{r(x)>\frac{\pi}{2}}\int_{|y-x|<1}\sigma_{R}(y)\,dy$, then $\exists x_{\pi,R}\in\mathbb{R}^{3}$ with $r\left(x_{\pi,R}\right)>\frac{\pi}{2}$, such that
$\int_{\left|y-x_{\pi,R}\right|<1}\sigma_{R}(y)\,dy>\epsilon_{\pi,R}-\frac{1}{\pi}$
On the other hand, due to the axial symmetry of $\sigma_{R}(y),\sigma_{R}(y)$ has mass $\geq{cr}\left(x_{\pi,R}\right)\left(\epsilon_{\pi,R}-\frac{1}{\pi}\right)$ in the torus obtained by revolving the ball $\left|y-x_{\pi,R}\right|<1$ around the z-axis, which implies $\epsilon_{\pi,R}\leq\frac{m}{{cr}\left(x_{\pi,R}\right)}+\frac{1}{\pi}<\frac{2m}{c\pi}+\frac{1}{\pi}$. By choosing $c$ sufficiently large, we have $\epsilon_{\pi,R}<\frac{c}{\pi}$.
For all $x$ with $r(x)>\pi$, notice due to $r<\frac{1+\pi}{2}$, we can choose finite balls centered at $x_{n}$ of radius 1 with $r\left(x_{n}\right)>\frac{\pi}{2}$ covering 1<|yx|<r1<|y-x|<r, and the number of balls can be $\leq cr^{3}$. Similar to Lemma 2.24, together with $\epsilon_{\pi,R}<\frac{c}{\pi}$, we have
$V_{\sigma_{R}}(x)$ $=\int_{|y-x|<1}\frac{\sigma_{R}(y)}{|y-x|}\,dy+\int_{1<|y-x|<r}\frac{\sigma_{R}(y)}{|y-x|}\,dy+\int_{|y-x|>r}\frac{\sigma_{R}(y)}{|y-x|}\,dy$